An education rehearser claims that at most 5% of working college students are employed as teachers or teaching assistants. In a random sample of 200 working college students, 12 of them are employed as teachers or teaching assistants. Is there enough evidence to support your thinking at α = 0.05?

1. What is the p-value
Answer #1

Solution:

Level of significance = \alpha = 0.05

Point estimate = sample proportion = р = x / n = 0.06

This a right (One) tailed test.

The null and alternative hypothesis is,

Ho: p \leq 0.05

Ha: p > 0.05

Test statistics

z = (р - Po ) / \sqrt{} Po*(1-Po) / n

= ( 0.06 - 0.05) / \sqrt{} (0.05*0.95) /200

= 0.649

P-value = P(Z > z)

= 1 - P( Z < 0.649 )

= 1 - 0.7418

= 0.2582

The P-value is 0.2582.